A television screen is 16cm by 22cm. The manufacturer wants to increase the outer edge by a uniform width to double the area. Find the uniform width to the nearest cm.
Answer by ikleyn(52867)

.

New dimensions are  (16+2x) by (22+2x).


The area equation is

    (16+2x) * (22+2x) = 2*16*22


Simplify and solve for x

    16*22 + 44x + 32x + 4x^2 = 2*16*22

    4x^2 + 76x - 16*22 = 0

     x^2 + 19x - 88 = 0


At this point, use the quadratic formula and choose positive root.


ANSWER.  3.851 cm (approximately).

Solved.

--------------

If you want to see many other similar solved problems, look into the lessons
    - Problems on the area and the dimensions of a rectangle surrounded by a strip
    - Cynthia Besch wants to buy a rug for a room
    - Problems on a circular pool and a walkway around it
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic
"Dimensions and the area of rectangles and circles and their elements".

Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.